Determine the constant that should be added to the binomial so that it becomes a perfect square trinomial. Then write and factor the trinomial.
x^2-4x
a)16;x^2-4x+16=(x-4)^2
b)-4;x^2-4x-4=(x-2)^2
c)-16;x^2-4x-16=(x-4)^2
d)4;x^2-4x+4=(x-2)^2

Answer by edjones(8007)

x^2-4x
divide the 2nd term by 2 and square the result.
-4/2=-2
-2^2=4
x^2-4x+4 Ans.
.
Ed